STABLE C∗-ALGEBRAS

MIKAEL RØRDAM

Abstract. We give a survey of known results, and add some new results, about stable C^{∗}-algebras. The main themes in the article are to characterize stable C^{∗}- algebras, to discuss permanence properties of the class of stable C^{∗}-algebras, and to investigate the relation between this topic and the structure problem for simple C^{∗}-algebras.